OpenStudy (anonymous):

the fundtion f is one to one. find its inverse: f(x)=cube root (x+3) please show me step by step:). thanks....

OpenStudy (anonymous):

x = cube root (y+3) x^3 = y+3 y = x^3-3

OpenStudy (anonymous):

y=cubert(x+3) swap the y and the x variable x=cubert(y+3) cube either side (so you can get rid of that cube root) x^3=[cubert(y+3)]^3 the 3rd power and cube root cancel out x^3=y+3 solve for y y=x^3-3